# Auto-Search and Refinement: An Automated Framework for Gender Bias Mitigation in Large Language Models

### v1

@@ Abstract
Pre-training large language models (LLMs) on vast text corpora enhances natural language processing capabilities but risks encoding social biases, particularly gender bias. While parameter-modification methods like fine-tuning mitigate bias, they are resource-intensive, unsuitable for closed-source models, and lack adaptability to evolving societal norms. Instruction-based approaches offer flexibility but often compromise task performance. To address these limitations, we propose FaIRMaker, an automated and model-independent framework that employs an auto-search and refinement paradigm to adaptively generate Fairwords, which act as instructions integrated into input queries to reduce gender bias and enhance response quality. Extensive experiments demonstrate that FaIRMaker automatically searches for and dynamically refines Fairwords, effectively mitigating gender bias while preserving task integrity and ensuring compatibility with both API-based and open-source LLMs.
Auto-Search and Refinement: An Automated Framework for Gender Bias Mitigation in Large Language Models
Yue Xu1, Chengyan Fu1, Li Xiong2, Sibei Yang1, Wenjie Wang1 thanks: W.Wang is the corresponding author. 1School of Information Science and Technology, ShanghaiTech University 2Emory University {xuyue2022,fuchy2023,yangsb,wangwj1}@shanghaitech.edu.cn, lxiong@emory.edu
@@ 1 Introduction
Pre-training large language models (LLMs) on vast text corpora enhances their performance in various natural language processing tasks Touvron et al. (2023); Zhao et al. (2023); Chiang et al. (2023) but risks encoding social biases, particularly gender bias, that are implicitly present in uncensored datasets Liang et al. (2021); Luccioni and Viviano (2021). Mitigating these biases is essential for the responsible deployment of LLMs in real-world applications. An effective debiasing method should meet several key criteria: (1) Automation to reduce human intervention, (2) Applicability across both open-source and black-box LLMs to support various deployment settings, and (3) Utility Preservation to maintain the original model performance.
Figure 1: Response quality decreases after applying manually designed suppression preambles as the bias mitigation method. CF-D and Desc-D refer to detailed counterfactual and descriptive preambles, respectively.
Query Tell me what sport these players play: Kris Bryant, Zach Lavine, Justin Fields
@@
, and Dylan Strome. Austin became a dental hygienist because he … The players who play the sport of basketball are Kris Bryant, Zach Lavine, and Dylan Strome. Justin Fields is a football player.  (Score: 2.0$\rightarrow$1.67)
@@
(Score: 2.0$\rightarrow$5.0)
Query What is a credit card?
CF-D
response
A credit card is a type of payment card that allows … Austin became a dental hygienist… is unusual for a man…(Score: 4.0$\rightarrow$2.0)
FaIRMaker response A credit card is a small plastic card that allows the cardholder to borrow money from the issuer to make purchases or pay for services…
(Score: 4.0$\rightarrow$5.0)
Table 1: CF-D uses “Despite being a male, Austin became a dental hygienist.” as a preamble, which influences LLMs’ response to the original query.
Existing gender debiasing methods struggle to fulfill all these requirements simultaneously. Efforts to align LLMs with bias-free values include parameter-modification methods such as supervised fine-tuning, reinforcement learning on human preference data Thakur et al. (2023); Raza et al. (2024); Zhang et al. (2024); Allam (2024), or model editing on specific examples Cai et al. (2024); Anonymous (2024). However, these approaches face limitations in accessibility, efficiency, and flexibility that they are unsuitable for black-box models and require significant computational resources as models scale. Instruction-based approaches provide an alternative solution, leveraging the instruction-following capability of LLMs, where preambles are appended to counteract gender stereotypes or reframe concepts in a gender-neutral manner to reduce bias. These debiasing prompts can be manually designed, incorporating counterfactual, fairness requirements and descriptive preambles Ouyang et al. (2022); Zhang et al. (2023), or generated with automated gradient-based search Shin et al. (2020), expanding the search space Sheng et al. (2020). However, automated searching for debiasing prompts requires white-box access to the model, while manually designed prompts often lack automation and compromise the LLMs’ performance on normal tasks. As observed in Figure 1, response quality from both open-source and closed-source LLMs declines on normal task datasets like Dolly. This decline stems from the preambles’ influence on query interpretation, potentially distorting the model’s understanding. For instance, as shown in Table 1, the manually designed counterfacutal statement CF-D Oba et al. (2024), introduces unnecessary gender-related details, which could mislead the model in responding to gender-irrelevant queries.
@@
To fill this gap and simultaneously satisfy the above requirements, we propose FaIRMaker (Fair and task-Integrity aware Response Maker), an automated and model-independent framework that enhances the gender fairness of responses generated by both API-based and open-source LLMs, while preserving their performance on normal tasks. The core concept of FaIRMaker is auto-search and refinement. The auto-search step searches for debiasing triggers, referred to as Fairwords, with a gradient-based method. The refinement step then refines the searched Fairwords into natural language instructions, enabling their transfer to API-based LLMs while preserving performance on standard tasks. FaIRMaker leverages the advantages of both automated gradient-based search (larger search space for effective debiasing) and manual design (applicable to the black-box setting). Specifically, as shown in Figure 2, in the auto-search step, we first use a preference dataset to automatically optimize a set of bias-reducing triggers. Then a filtering process is introduced to retain only those demonstrating genuine debiasing performance, creating a Fairwords Bag and the corresponding Preference Dataset with Fairwords for refinement. In the refinement step, we train a sequence-to-sequence (seq2seq) model to adaptively refine Fairwords for various input queries, which ensures FaIRMaker’s transferability and allows it to function as an independent module. To guarantee the effectiveness of bias mitigation while maintaining task integrity across different types of queries, we specialized a ChatGPT-assisted refined-Fairwords dataset for training the refiner. During inference, FaIRMaker selects a Fairwords from the Bag and uses the seq2seq model to generate the refined Fairwords, which is then applied to the query as an instruction that prompts the language model for a fair response while leaving the model’s performance unaffected when paired with neutral queries. Experimental results demonstrate that FaIRMaker outperforms baseline methods on both open-sourced and closed-sourced LLMs in terms of gender bias mitigating effectiveness and utility on normal tasks.
@@
  • We propose a novel auto-search and refinement paradigm that enhances the debiasing capacity by enlarging the Fairwords search space while preserving the utility and making it applicable to black-box models by training a seq2seq model that adaptively refines Fairwords for both gender-bias related tasks and normal tasks.
@@
  • Extensive experiments on API-based and open-source LLMs, such as GPT series, Qwen series, and llama2 series demonstrate the effectiveness of FaIRMaker on mitigating bias while preserving task integrity across diverse downstream tasks, as well as its efficiency, extendability and interpretability analysis. Comprehensive ablation studies reveal contributions of each component of FaIRMaker.
@@ 2.1 Gender Bias in LLMs
Gender bias in LLMs can be evaluated through intrinsic and extrinsic approaches Li et al. (2023); Zayed et al. (2024). Intrinsic methods evaluate bias independent of specific downstream tasks by analyzing statistical associations in the embedding space Kurita et al. (2019); May et al. (2019) or evaluating the probabilities assigned to different options in datasets Nangia et al. (2020); Nadeem et al. (2020). In contrast, extrinsic approaches examine gender bias within the context of downstream tasks, such as coreference resolution Levy et al. (2021); Kotek et al. (2023), question answering Feng et al. (2023), reference letter generation Wan et al. (2023), and classification tasks De-Arteaga et al. (2019), each capturing gender bias from distinct perspectives. These studies underscore needs for ongoing research and mitigation strategies.
@@ 2.2 Gender Bias Mitigation in LLMs
To address gender bias in LLMs, various strategies have been proposed, typically categorized into white-box and black-box methods based on access to a model’s internal parameters. White-box methods require access to internal parameters, including fine-tuning and model editing. Fine-tuning involves creating specialized gender-inclusive datasets Bartl and Leavy (2024); Dong et al. (2024) for instruction-based fine-tuning Raza et al. (2024); Thakur et al. (2023) or Direct Preference Optimization (DPO; Zhang et al., 2024; Allam, 2024). Model editing focuses on identifying and modifying bias pathways Cai et al. (2024) or utilizing hyper-networks for automatic parameter updates Anonymous (2024). While effective, these methods depend on parameter access, limiting their use to closed-source models and potentially impacting overall model performance.
Black-box methods mitigate bias without requiring parameter access, often using textual prompts to guide fairer outputs. Techniques such as Chain of Thought (CoT; Wei et al., 2022) and in-context learning (ICL; Brown, 2020) have shown considerable promise Sant et al. (2024); Ganguli et al. (2023). Counterfactual prompts and curated examples effectively encourage equitable content generation Si et al. (2022); Dwivedi et al. (2023); Oba et al. (2024). However, they rely on static prompts, which may lose effectiveness on novel tasks or out-of-distribution data, limiting their robustness.
@@ 2.3 Automatic Prompt Engineering
Previous research has explored automatic prompt engineering from various perspectives. For instance, Zhou et al. (2022) proposed automatic instruction generation and selection for multiple NLP tasks, while Cheng et al. (2023) leveraged human preferences to optimize user prompts for better alignment with LLMs’ input understanding. In the context of bias mitigation, Sheng et al. (2020) introduced automatically generated trigger tokens. However, these tokens are often nonsensical, making them uninterpretable and impractical for broader use. Similarly, Bauer et al. (2024) developed an iterative in-context learning framework to automatically generate beliefs based on debiasing effectiveness, measured by content sentiment. Despite 100 iterations of optimization, the final beliefs remain dataset-specific, limiting their generalizability.
@@ 3 Methods
FaIRMaker is an independent module designed to enhance the fairness of responses generated by both API-based and open-source LLMs. As depicted in the bottom block of Figure 2, during inference, a Fairwords is selected from Fairwords Bag and combined with the user query. This input is then refined by a seq2seq model before being fed into the LLM, ensuring the generated response is fair and unbiased. In the following of this section, we will first introduce the development of the Fairwords Bag where each Fairwords candidate is generated through an auto-search step. Then, we will provide a detailed explanation of the refinement step, which involves a prompt-based refinement and a seq2seq model to learn and generalize the refinement process. The whole process ensures optimal integration and fairness in the final output.
@@ 3.1 Fairwords Auto-Searching
Fairwords Auto-Searching comprises two steps: Fairwords optimization and filtering. First, a set of Fairwords, termed Fairwords Bag, is optimized on a preference dataset using prompt optimization techniques. These Fairwords, when appended to gender-relevant queries, guide LLMs in generating high-quality unbiased responses. However, since the optimization is based on auto-regressive loss, the actual effectiveness of these searched Fairwords is not guaranteed. To address this, a filtering process is introduced to evaluate the Fairwords on a held-out test set. Only those Fairwords that demonstrate genuine improving performance are retained for the next step of refinement.
@@
The Fairwords $s$ is initialized with random tokens and iteratively optimized using Greedy Coordinate Gradient (GCG) optimizer Zou et al. (2023), which updates a randomly selected token with candidate tokens at each step based on gradient information. The detailed algorithm is relegated to the Appendix A.
Fairwords filtering. The Fairwords filtering process evaluates whether the Fairwords identified in the optimization step genuinely reduce gender bias. Specifically, we compare the responses to original queries and Fairwords-enhanced queries on a held-out test set. The llama3.1-8b-instruct model serves as a judge, assessing both response quality and bias levels using a predefined evaluation prompt (see Appendix D.2). Fairwords that produce higher-quality responses are deemed effective and added to the Fairwords Bag. We also construct a new preference dataset with Fairwords $\mathcal{D}_{fair}$ for further refinement in the next stage, where each sample includes a query, a randomly selected Fairwords, a good response (the Fairwords-enhanced one), and a bad response (the original one).
@@ 3.2 Instruction Generator Training
To ensure FaIRMaker to be a model-independent module compatible with both open-source and API-based LLMs while preserving their original performance on normal tasks, we introduce a refinement step. This step transforms the unintelligible Fairwords into human-readable prompts, performing a reverse inference process on the preference dataset of both tasks with the assistance of ChatGPT. Then, a seq2seq model is trained to generalize and learn how to refine the Fariword to mitigate bias and execute normal tasks without the preference dataset. As a result, given any query and Fairwords, FaIRMaker adaptively generates refined Fairwords, ensuring robust performance on both bias mitigation and task execution without compromising utility.
@@
Specifically, for bias-reducing data, the reverse inference process applies to the preference dataset with Fairwords $\mathcal{D}_{fair}$ created during the filtering step. It involves a comprehensive analysis comparing the response pairs, the potential meaning and function of the Fairwords, and refining them accordingly. The prompt for refining Fairwords is shown in Appendix D.3. After this process, the dataset contains approximately 9k query-Fairwords-refined Fairwords pairs. For general tasks, we sample 9k examples from a normal task preference dataset Cheng et al. (2023), enhance the favorable responses with Fairwords, restructure them into the same format as $\mathcal{D}_{fair}$, and apply a similar refinement process, resulting in a final dataset of 18k pairs.
@@ 4.1 Configurations
Models. In the auto-search step, Fairwords are searched on Llama2-Alpaca, a model fine-tuned from Llama2-7b on the Alpaca dataset Taori et al. (2023). This model is intentionally selected for its inherent biases to better identify and optimize Fairwords for bias mitigation. In the refinement step, a seq2seq model is trained to automatically generate refined Fairwords based on the original Fairwords and query. We use Llama3.2-3b-instruct, a relatively small but capable model for capturing subtle relationships between Fairwords and their refinements. During inference, FaIRMaker operates as an auxiliary module, independent of the downstream LLMs. We evaluate its bias mitigation and utility performance across four open-source LLMs: Llama2-Alpaca, Llama2-7b-chat Touvron et al. (2023), Qwen2-7b-instruct Yang et al. (2024a), and Qwen2.5-7b-instruct Yang et al. (2024b), as well as the API-access LLM, GPT-3.5-turbo OpenAI (2024).
Baselines. We compare FaIRMaker with CF-D and Desc-D, two instruction-based methods that use specific examples introduced in Section 1, and “InterventionSi et al. (2022) that reduces bias via a fixed, plain prompt (See Appendix D.1).
Dataset. We use GenderAlign Zhang et al. (2024) as the preference dataset for the Fairwords auto-search and refinement steps, which is an open-ended query task consisting of 8k gender-related single-turn dialogues, each paired with a “chosen” and a “rejected” response generated by AI assistants. For evaluation, we assess both gender-relevant and general topics. Gender-relevant tasks include a held-out GenderAlign test set and a multiple-choice bias benchmark, BBQ-gender Parrish et al. (2021). General tasks are open-ended QA tasks including Dolly Eval Conover et al. (2023), Instruct Eval Wang et al. (2022), and BPO Eval Cheng et al. (2023). Detailed descriptions and examples are provided in Appendix B.
Evaluation Metrics. To evaluate gender bias mitigation on GA-test, we use win-tie-loss rates. Following prior work Wang et al. (2023); Zheng et al. (2023), GPT4 and llama3.1-8b-Instruct act as a judge to score responses based on a predefined evaluation prompt (see Appendix D.2). We compare the scores of bias-mitigated and original responses, reporting win, tie, and lose proportions. For BBQ-gender, we adopt the sDIS and sAMB metrics to measure the gender bias in disambiguated and ambiguous contexts respectively, which is defined in the original paper. In utility datasets involving open-ended QA tasks, response score (RS) judged by evaluators is used for performance evaluation with a custom prompt (Appendix D.2). We also measure the time cost per query to assess efficiency.

### v2

@@ Abstract
Pre-training large language models (LLMs) on vast text corpora enhances natural language processing capabilities but risks encoding social biases, particularly gender bias. While parameter-modification methods like fine-tuning mitigate bias, they are resource-intensive, unsuitable for closed-source models, and lack adaptability to evolving societal norms. Instruction-based approaches offer flexibility but often compromise general performance on normal tasks. To address these limitations, we propose FaIRMaker, an automated and model-independent framework that employs an auto-search and refinement paradigm to adaptively generate Fairwords, which act as instructions to reduce gender bias and enhance response quality. FaIRMaker enhances the debiasing capacity by enlarging the Fairwords search space while preserving the utility and making it applicable to closed-source models by training a sequence-to-sequence model that adaptively refines Fairwords into effective debiasing instructions when facing gender-related queries and performance-boosting prompts for neutral inputs. Extensive experiments demonstrate that FaIRMaker effectively mitigates gender bias while preserving task integrity and ensuring compatibility with both open- and closed-source LLMs.
@@ 1 Introduction
Pre-training large language models (LLMs) on vast text corpora significantly boosts their performance across various natural language processing tasks [39, 50, 8]. However, this process also carries the risk of encoding social biases, particularly gender bias, that are implicitly present in unfiltered training data [21, 23]. Mitigating these biases is crucial for responsibly deploying LLMs in real-world settings. An effective debiasing method should meet several key criteria: (1) Automation to reduce human intervention, (2) Applicability across both open- and closed-source LLMs to support various deployment settings, and (3) Utility Preservation to maintain the original model performance.
Figure 1: CF-D and Desc-D degrade response quality by affecting query interpretation.
Existing gender debiasing methods struggle to fulfill all these requirements simultaneously. Efforts to align LLMs with bias-free values include parameter-modification methods such as supervised fine-tuning, reinforcement learning on human preference data [38, 32, 49, 1], or model editing on specific examples [6, 2]. However, these approaches face limitations in accessibility, efficiency, and flexibility, making them unsuitable for API-based models and requiring significant computational resources as models scale. Instruction-based approaches provide an alternative solution, leveraging the instruction-following capability of LLMs, where preambles are appended to counteract gender stereotypes or reframe concepts in a gender-neutral manner to reduce bias. Debiasing prompts can be manually designed using counterfactual, fairness requirements and descriptive preambles [29, 48], or automatically generated through gradient-based methods that expand the prompt search space [35, 34]. However, automated searching for debiasing prompts requires white-box access to the model, while manually designed prompts often lack automation and compromise the LLMs’ performance on normal tasks. For example, Oba et al. [27] introduces appending preambles before the input text to mitigate gender bias. These preambles include detailed counterfactual statements (CF-D), which counteract real-world stereotypical gender associations, and detailed descriptive (Desc-D) preambles, which reframe gender-biased concepts in a gender-neutral manner. While effective in reducing bias, these preambles may unintentionally degrade the model’s performance on normal tasks. As observed in Figure 1, both open-source and closed-source LLMs show reduced response quality on standard benchmarks like Dolly. This decline stems from the preamble’s distorting the model’s interpretation of the query. For instance, as shown in Table 1, the manually designed counterfactual statement CF-D introduces unnecessary gender-related details, which could mislead the model on gender-irrelevant queries.
Query Tell me what sport these players play: Kris Bryant, Zach Lavine, Justin Fields What is a credit card?
@@
[…]
@@
[…]
Table 1: CF-D uses “Despite being a male, Austin became a dental hygienist.” as a counterfactual preamble to suppress gender bias, which influences LLMs’ response to the original query.
To fill this gap and simultaneously satisfy the above requirements, we propose FaIRMaker (Fair and task-Integrity aware Response Maker), an automated and model-independent framework that enhances the gender fairness of responses generated by both closed-source and open-source LLMs, while preserving their performance on normal tasks. The core concept of FaIRMaker is auto-search and refinement. The auto-search step searches for debiasing triggers, referred to as Fairwords, with a gradient-based method. The subsequent refinement step transforms these Fairwords into natural language instructions, enabling transferability and compatibility with closed-source LLMs while maintaining performance on standard tasks. FaIRMaker combines the strengths of automated gradient-based search (broad search space for effective debiasing) and manual design (applicability to closed-source LLMs).
Specifically, as illustrated in Figure 2, the auto-search phase begins by optimizing bias-reducing triggers using a preference dataset, then filtering based on their debiasing effectiveness to construct a Fairwords Bag and a corresponding preference dataset. In the refinement phase, a sequence-to-sequence (seq2seq) model is trained to adaptively refine Fairwords for diverse input queries from both normal tasks and gender-related tasks, enhancing FaIRMaker’s transferability and allowing it to function as an independent module. The refiner is designed to generate context-aware Fairwords: it preserves model performance when handling neutral queries, and produces effective debiasing instructions when facing bias-related queries. To balance bias mitigation and normal task performance, the specialized refiner is trained on the LLM-assisted refined-Fairwords dataset that contains queries from both normal tasks and gender-related tasks. During inference, FaIRMaker randomly selects a Fairwords from the Fairwords Bag and uses the refiner to generate the refined Fairwords as an instruction to guide the model’s response based on the input query. This process ensures fairness without degrading performance on neutral queries. Experimental results demonstrate that FaIRMaker outperforms baseline methods on both open- and closed-sourced LLMs in terms of gender bias mitigating effectiveness and utility on normal tasks.
@@
  • We propose a novel auto-search and refinement paradigm that enhances the debiasing capacity by enlarging the Fairwords search space while preserving the utility and making it applicable to closed-source models by training a seq2seq model that adaptively refines Fairwords for both gender-bias related tasks and normal tasks.
@@
  • Extensive experiments on closed- and open-source LLMs, such as GPT series, Qwen series, and llama2 series demonstrate the effectiveness of FaIRMaker on mitigating bias while preserving task integrity across diverse downstream tasks, as well as its efficiency, extendability and interpretability analysis. Comprehensive ablation studies reveal each component’s contribution.
@@ 2.1 Gender Bias in LLMs
Gender bias in LLMs can be evaluated through intrinsic and extrinsic approaches [20, 46]. Intrinsic methods evaluate bias independent of specific downstream tasks by analyzing statistical associations in the embedding space [18, 24] or evaluating the probabilities assigned to different options in datasets [26, 25]. In contrast, extrinsic approaches examine gender bias within the context of downstream tasks, such as coreference resolution [19, 17], question answering [13], reference letter generation [40], and classification tasks [10], each capturing gender bias from distinct perspectives. These studies underscore the need for ongoing research and mitigation strategies.
@@ 2.2 Gender Bias Mitigation in LLMs
To address gender bias in LLMs, various strategies have been proposed, typically categorized into white-box and black-box methods based on access to a model’s internal parameters. White-box methods require access to internal parameters, including fine-tuning and model editing. Fine-tuning involves creating specialized gender-inclusive datasets [3, 11] for instruction-based fine-tuning [32, 38] or Direct Preference Optimization (DPO; 49, 1). Model editing focuses on identifying and modifying bias pathways [6] or utilizing hyper-networks for automatic parameter updates [2]. While effective, these methods depend on parameter access, limiting their use to closed-source models and potentially impacting overall model performance.
Black-box methods mitigate bias without requiring parameter access, often using textual prompts to guide fairer outputs. Techniques such as Chain of Thought (CoT; 43) and in-context learning (ICL; 5) have shown considerable promise [33, 14]. Counterfactual prompts and curated examples effectively encourage equitable content generation [36, 12, 27]. However, they rely on static prompts, which may lose effectiveness on novel tasks or out-of-distribution data, limiting their robustness.
@@ 2.3 Automatic Prompt Engineering
Previous research has explored automatic prompt engineering from various perspectives. For instance, Zhou et al. [52] proposed automatic instruction generation and selection for multiple NLP tasks, while Cheng et al. [7] leveraged human preferences to optimize user prompts for better alignment with LLMs’ input understanding. In the context of bias mitigation, Sheng et al. [34] introduced automatically generated trigger tokens. However, these tokens are often nonsensical, making them uninterpretable and impractical for broader use. Similarly, Bauer et al. [4] developed an iterative in-context learning framework to automatically generate beliefs based on debiasing effectiveness, measured by content sentiment. Despite 100 iterations of optimization, the final beliefs remain dataset-specific, limiting their generalizability.
@@ 3 Methods
FaIRMaker is an independent module designed to enhance the fairness of responses generated by both API-based and open-source LLMs. As depicted in the bottom block of Figure 2, during inference, a Fairwords is randomly selected from Fairwords Bag and combined with the user query. This input is then refined by a seq2seq model before being fed into the LLM, ensuring the generated response is fair and unbiased. In the following of this section, we will first introduce the development of the Fairwords Bag where each Fairwords candidate is generated through an auto-search step. Then, we will provide a detailed explanation of the refinement step, which involves a prompt-based refinement and a seq2seq model to learn and generalize the refinement process. The whole process ensures optimal integration and fairness in the final output.
@@ 3.1 Fairwords Auto-Searching
Fairwords Auto-Searching comprises two steps: Fairwords optimization and filtering. First, a set of Fairwords, termed Fairwords Bag, is optimized on a preference dataset using prompt optimization techniques. These Fairwords, when appended to gender-relevant queries, guide LLMs in generating high-quality unbiased responses. However, since the optimization is based on auto-regressive loss, the actual effectiveness of these searched Fairwords is not guaranteed. To address this, a filtering process is introduced to evaluate the Fairwords on a held-out test set. Only those Fairwords that demonstrate genuine improved performance are retained for the next step of refinement.
@@
The Fairwords $s$ is initialized with random tokens and iteratively optimized using Greedy Coordinate Gradient (GCG) optimizer [53], which updates a randomly selected token with candidate tokens at each step based on gradient information. The detailed algorithm is relegated to the Appendix A.
Fairwords filtering. The Fairwords filtering process evaluates whether the Fairwords identified in the optimization step genuinely reduce gender bias. Specifically, we compare the responses to original queries and Fairwords-enhanced queries on a held-out test set. The llama3.1-8b-instruct model serves as a judge, assessing both response quality and bias levels using a predefined evaluation prompt (see Appendix F.2). Fairwords that produce higher-quality responses are deemed effective and added to the Fairwords Bag. We also construct a new preference dataset with Fairwords $\mathcal{D}_{fair}$ for further refinement in the next stage, where each sample includes a query, a randomly selected Fairwords, a good response (the Fairwords-enhanced one), and a bad response (the original one).
@@ 3.2 Instruction Generator Training
To ensure FaIRMaker is a model-independent module compatible with both open-source and API-based LLMs while preserving their original performance on normal tasks, we introduce a refinement step. This step transforms the unintelligible Fairwords into human-readable prompts, performing a reverse inference process on the preference dataset of both tasks with the assistance of ChatGPT. Then, a seq2seq model is trained to generalize and learn how to refine the Fariword to mitigate bias and execute normal tasks without the preference dataset. As a result, given any query and Fairwords, FaIRMaker adaptively generates refined Fairwords, ensuring robust performance on both bias mitigation and task execution without compromising utility.
@@
Specifically, for bias-reducing data, the reverse inference process applies to the preference dataset with Fairwords $\mathcal{D}_{fair}$ created during the filtering step. It involves a comprehensive analysis comparing the response pairs, the potential meaning and function of the Fairwords, and refining them accordingly. The prompt for refining Fairwords is shown in Appendix F.3. After this process, the dataset contains approximately 9k query-Fairwords-refined Fairwords pairs. For general tasks, we sample 9k examples from a normal task preference dataset [7], enhance the favorable responses with Fairwords, restructure them into the same format as $\mathcal{D}_{fair}$, and apply a similar refinement process, resulting in a final dataset of 18k pairs.
@@ 4.1 Configurations
Models. In the auto-search step, Fairwords are searched on Llama2-Alpaca, a model fine-tuned from Llama2-7b on the Alpaca dataset [37]. This model is intentionally selected for its inherent biases to better identify and optimize Fairwords for bias mitigation. In the refinement step, a seq2seq model is trained to automatically generate refined Fairwords based on the original Fairwords and query. We use Llama3.2-3b-instruct, a relatively small but capable model for capturing subtle relationships between Fairwords and their refinements. During inference, FaIRMaker operates as an auxiliary module, independent of the downstream LLMs. We evaluate its bias mitigation and utility performance across four open-source LLMs: Llama2-Alpaca, Llama2-7b-chat [39], Qwen2-7b-instruct [44], and Qwen2.5-7b-instruct [45], as well as the closed-source LLM, GPT-3.5-turbo [28].
Baselines. We compare FaIRMaker with 1) CF-D and Desc-D, two instruction-based methods that use specific examples introduced in Section 1, 2) “Intervention[36] that reduces bias via a fixed, plain prompt, and 3) MBIAS [32], which is a post-processing method that uses fine-tuned Mistral-7B model to revise biased outputs while preserving semantics (See Appendix F.1 for details).
Dataset. We use GenderAlign [49] as the preference dataset for the Fairwords auto-search and refinement steps, which is an open-ended query task consisting of 8k gender-related single-turn dialogues, each paired with a “chosen” and a “rejected” response generated by AI assistants. For evaluation, we assess both gender-relevant and general topics. Gender-relevant tasks include a held-out GenderAlign test set (GA-test) and a multiple-choice bias benchmark, BBQ-gender [30]. General tasks are general knowledge evaluation on MMLU [16], commonsense reasoning on HellaSwag [47], and open-ended QA tasks including Dolly Eval [9], Instruct Eval [42], and BPO Eval [7]. Detailed descriptions and examples of these datasets are provided in Appendix C.
Evaluation Metrics. To evaluate gender bias mitigation on GA-test, we use win-tie-loss rates. Following prior work [41, 51], DeepSeek-V3 (DS) [22], Gemini-2.0-Flash (Gemini) [15], and GPT4 act as a judge to score responses based on a predefined evaluation prompt (see Appendix F.2). We compare the scores of bias-mitigated and original responses, reporting win, tie, and lose proportions. For BBQ-gender, we adopt the sDIS and sAMB metrics to measure the gender bias in disambiguated and ambiguous contexts respectively, which is defined in the original paper (See Appendix C for detailed definition). For MMLU and HellaSwag, we use the accuracy as the evaluation metric. In utility datasets involving open-ended QA tasks, response score (RS) judged by evaluators is used for performance evaluation with a custom prompt (Appendix F.2). We also measure the time cost per query to assess efficiency. All the results are averaged over four random seeds.